The density of a gas is 0.28 1b/ft. What volume does 12.0 lbs of the gas occupy? What is the specific volume of the gas?
1
Expert's answer
2021-10-31T18:21:50-0400

Density"\rho=0.28lb/ft^3=4.48517kg/m^3"

"Mass=12lbs=5.44311kg"

"Volume (v)=\frac{m}{\rho}"

Put value

"v=\frac{5.44311}{4.48517}=1.213m^3"

Specific volume


"V'=\frac{v}{m}"

Where (m)=5.44311kg

"V'=\frac{1.214}{5.44311}=0.22m^3/kg"